OpenStudy (cenaida):

Solve the following system of equations using elimination: 3x + 5y = 4 4x + 2y = -18 A. (-12, 8) B. (-7, 5) C. (-4, -1) D. (-2, 2)

OpenStudy (cenaida):

so is it A?

OpenStudy (anonymous):

hold on i gave you the wrong instructions, let me do it for you

OpenStudy (cenaida):

ok

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

you were supposed to multiply the first one by -4 and the second one by 3 so your x's would be -12 and 12, so they'd cancel out. y would = 5, then if you plugged it in as 3x + 5(5) = 4, you'd solve and get x= -7

OpenStudy (anonymous):

3x + 5y = 4 -->(-2) 4x + 2y = -18 -->(5) -------------------- -6x - 10y =-8 (result of multiplying by -2) 20x + 10y = - 90 (result of multiplying by 5) ---------------------add 14x = -98 x = -7 3x + 5y = 4 3(-7) + 5y = 4 -21 + 5y = 4 5y = 4 + 21 5y = 25 y = 5 wowisuckatschool is correct...it is b

OpenStudy (cenaida):

Thanks you guys!